OpenStudy (howard-wolowitz):

Can someone help me here:

OpenStudy (phi):

when the exponent is 0, as in 2^0, what value is that ? any idea?

OpenStudy (howard-wolowitz):

I'm not sure what you mean, but do you mean that d because of that the graph it would be rising?

OpenStudy (phi):

They want you to know that 2^-1= 1/2 2^0 = ? 2^1 = 2 2^2 = 4 2^3 = 8 Have you learned this ?

OpenStudy (phi):

notice the pattern. to get from 2^3 to 2^2 you divide by 2 example 2^3 is 8. divide by 2: 8/2= 4 , and that is 2^2 if you keep doing that, you can figure out 2^0

OpenStudy (howard-wolowitz):

yes i have

OpenStudy (phi):

you should find 2^0 is 1 and try to remember that x^0 (for any number x except x being 0), you get x^0 = 1 now what is -3* 2^0 ? it is -3*1 in other words, when x is 0, y = -3 the curve goes through the point (0,-3) which curve does that ?

OpenStudy (howard-wolowitz):

top right one

OpenStudy (phi):

the top right shows the curve going through (0,3) is that what you want ?

OpenStudy (howard-wolowitz):



OpenStudy (phi):

oh, yes. top right looks good